Angela's father is 2 more than eight times her age. if the product of their ages is 300,how old is her father?
Answer by lwsshak3(11628)

Angela's father is 2 more than eight times her age. if the product of their ages is 300,how old is her father?
***
let x=angela's age
8x+2=father's age
..
x(8x+2)=300
8x^2+2x=300
8x^2+2x-300=0
(8x+50)(x-6)=0
8x+50=0
x=-50/8 (reject)
or
x-6=0
x=6
8x+2=48+2=50
how old is her father? 50